OpenStudy (anonymous):

What is the slope of the line between (3, –4) and (–2, 1)?

OpenStudy (cwrw238):

slope = 1-(-4) / -2-3 = 5 / -5 = -1

OpenStudy (agreene):

\[m=\frac{\Delta y}{\Delta x}\implies\frac{y_1-y_2}{x_1-x_2}\] (-4-1)/(3-(-2))=-1